### v1

@@ 2.2 Grid and coordinates
Compact objects are naturally described in spherical coordinates. Using body-fitted coordinate grids simplifies the inclusion of system symmetries and boundary conditions. Another degree of freedom is the usage of linear or non-uniform grid spacings. A common feature in similar codes is to allow uniform grid spacings in $\log r$ (logarithmic coordinate) and $-\cos\theta$ (equal area coordinate). The first allows for higher resolution close to the central object, while the second enforces constant charge density of each macro-particle in the meridional direction (Belyaev 2015a). These two properties have shown a reduction of numerical noise on the axis and increased code stability. Therefore, the OSIRIS-GR module has these two options available. Linear grid spacing options are also available, although these are not explored in the current paper. Another possibility available is to adopt logical coordinates, i.e. ($\tilde{r}$,$\tilde{\theta}$)=($\log r$,$-\cos\theta$), which possesses the two properties described above and retains uniform grid spacings.The background metric in the new coordinate system takes the form:
@@
$\tilde{A}^{\hat{r}}=\sqrt{\@tens{\gamma}_{\tilde{r}\tilde{r}}}\tilde{A}^{r}=\frac{\exp{\left(\tilde{r}\right)}}{\alpha}\exp{\left(-\tilde{r}\right)}A^{r}=\frac{1}{\alpha}A^{r}=\frac{1}{r}\frac{A^{\hat{r}}}{\alpha^{-1}}=A^{\hat{r}}.$ (4)
@@ 2.3 Maxwell’s equations
In the 3+1 formalism, Maxwell’s equations take the form (Komissarov 2011):
${E}=4\pi\rho,$ (5)
${B}=0,$ (6)
×${\nabla}\times\left(\alpha\@vec{E}+\@vec{\beta}\times\@vec{B}\right)=-\frac{\partial\@vec{B}}{\partial t},$ (7)
×${\nabla}\times\left(\alpha\@vec{B}-\@vec{\beta}\times\@vec{E}\right)=\frac{\partial\@vec{E}}{\partial t}+4\pi\left(\alpha\@vec{j}-\rho\@vec{\beta}\right),$ (8)
where ${E}$, ${B}$, ${j}$ and $\rho$ are physical quantities measured by fiducial observers in co-rotation with the absolute space (ZAMOs). In spherical coordinates, Maxwell’s equations are better described in the integral form using the Kelvin–Stokes theorem on a Yee-lattice (Belyaev 2015b; Cerutti et al. 2015). The OSIRIS-GR module adopts the spherical cell shown in Fig. 1.
@@
(×${\tilde{E}}\right)^{\hat{r}}_{i,j+1/2}=\frac{\alpha_{i}}{A^{\hat{r}}_{i,j+1/2}}\left(l^{\hat{\phi}}_{i,j+1}E^{\hat{\phi}}_{i,j+1}-l^{\hat{\phi}}_{i,j}E^{\hat{\phi}}_{i,j}\right),$ (9)
(×${\tilde{E}}\right)^{\hat{\theta}}_{i+1/2,j}=\frac{1}{A^{\hat{\theta}}_{i+1/2,j}}\left(l^{\hat{\phi}}_{i,j}\alpha_{i}E^{\hat{\phi}}_{i,j}-l^{\hat{\phi}}_{i+1,j}\alpha_{i+1}E^{\hat{\phi}}_{i+1,j}\right),$ (10)
(×${\tilde{E}}\right)^{\hat{\phi}}_{i+1/2,j+1/2}=\frac{l^{\hat{r}}_{i+1/2}}{A^{\hat{\phi}}_{i+1/2,j+1/2}}\left(\alpha_{i+1/2}\left(E^{\hat{r}}_{i+1/2,j}-E^{\hat{r}}_{i+1/2,j+1}\right)+\right.$
@@
(×${\tilde{B}}\right)^{\hat{r}}_{i+1/2,j}=\frac{\alpha_{i+1/2}}{A^{\hat{r}}_{i+1/2,j}}\left(l^{\hat{\phi}}_{i+1/2,j+1/2}B^{\hat{\phi}}_{i+1/2,j+1/2}-\right.$
@@
(×${\tilde{B}}\right)^{\hat{\theta}}_{i,j+1/2}=\frac{1}{A^{\hat{\theta}}_{i,j+1/2}}\left(l^{\hat{\phi}}_{i-1/2,j+1/2}\alpha_{i-1/2}B^{\hat{\phi}}_{i-1/2,j+1/2}-\right.$
@@
(×${\tilde{B}}\right)^{\hat{\phi}}_{i,j}=\frac{l^{\hat{r}}_{i}}{A^{\hat{\phi}}_{i,j}}\left(\alpha_{i}\left(B^{\hat{r}}_{i,j-1/2}-B^{\hat{r}}_{i,j+1/2}\right)+\right.$
@@
with ${\tilde{E}}\equiv\alpha\@vec{E}+\@vec{\beta}\times\@vec{B}$ and ${\tilde{B}}\equiv\alpha\@vec{B}-\@vec{\beta}\times\@vec{E}$, respectively. Expressions for the line and area elements are detailed in Appendix A. Up to this point, we have assumed $\beta^{\hat{\phi}}$ as the only non-zero component, with a purely radial dependence in $\alpha$, and neglected all azimuthal derivatives (axisymmetry). The evolution equations are then expressed as in the typical Yee (1966) scheme
${B}^{n+1/2}=\@vec{B}^{n}-\frac{\Delta t}{2}\left(\@vec{\nabla}\times\@vec{\tilde{E}}\right)^{n},$ (15)
${E}^{n+1}=\@vec{E}^{n}+\Delta t\left(\@vec{\nabla}\times\@vec{\tilde{B}}\right)^{n+1/2}-4\pi\Delta tJ^{n+1/2},$ (16)
${B}^{n+1}=\@vec{B}^{n+1/2}-\frac{\Delta t}{2}\left(\@vec{\nabla}\times\@vec{\tilde{E}}\right)^{n+1},$ (17)
with $J\equiv\alpha\@vec{j}-\rho\@vec{\beta}$ and $n$ being the temporal index. Equation (16) also shows that to correctly push in time the electric field we need to gather the particle’s half-step charge density and current (see Sect. 2.5 and Appendix B). Due to the complex coupling introduced by equations (11) and (14), the poloidal components are pushed in time before the azimuthal ones in equations (16) and (17). By applying the Newtonian limit, we retrieve the Minkowski version of the evolution equations in Belyaev (2015b); Cerutti et al. (2015). Notice that (9)-(14) can be applied to any axisymmetric system given a diagonal spatial metric (e.g. spherical, cylindrical, toroidal).
@@ 2.4 Particle pusher
In the 3+1 formalism, the equations of motion for a charged particle (Thorne & MacDonald 1982; Philippov et al. 2015a; Parfrey et al. 2019) are given by
$\frac{\mathrm{d}{r^{i}}}{\mathrm{d}t}\@vec{e}_{i}=\frac{\alpha}{\Gamma}\@vec{u}-\@vec{\beta},$ (18)
$\frac{\mathrm{d}{u^{\hat{i}}}}{\mathrm{d}t}\@vec{e}_{\hat{i}}=\frac{\alpha q}{m}\left(\@vec{E}+\frac{\@vec{u}}{\Gamma}\times\@vec{B}\right)+\alpha\Gamma\@vec{g}+\alpha\@vec{H}\cdot\@vec{u}+\@vec{F}_{RR}-\@vec{F}_{Coord},$ (19)
where ${g}$ is the gravitational acceleration, ${H}$ is the gravitomagnetic tensor, ${p}=m\Gamma\@vec{v}=m\@vec{u}$ is the FIDO-measured momentum, $q$ and $m$ are the charge and mass of each macro-particle, $\Gamma=\sqrt{\varepsilon+\@vec{u}\cdot\@vec{u}}$ is the particle’s special relativistic Lorentz factor, and $\varepsilon=0\;\left(=1\right)$ for massless (massive) particles. Equation (18) shows that the particle velocity is composed of the FIDO-measured velocity, $\alpha\@vec{u}/\Gamma$, and the FIDO’s velocity, ${\beta}$. Equation (19) details which forces dictate the particle dynamics. From left to right, we identify the Lorentz, the gravitational, the gravitomagnetic, the classical radiation reaction (Vranic et al. 2016; Landau & Lifschitz 1975), and the ”coordinate” forces. The last term comprises the generalized fictitious forces associated with the use of the orthonormal vector basis:
${F}_{Coord}=u^{\hat{i}}\frac{\mathrm{d}{\@vec{e}_{\hat{i}}}}{\mathrm{d}t}=\frac{\sqrt{\gamma_{ii}}}{\sqrt{\gamma_{kk}}}\Gamma^{i}_{kj}u^{\hat{k}}\frac{\mathrm{d}x^{j}}{\mathrm{d}t}\@vec{e}_{\hat{i}}=\Gamma^{\hat{i}}_{\hat{k}j}u^{\hat{k}}\frac{\mathrm{d}x^{j}}{\mathrm{d}t}\@vec{e}_{\hat{i}},\;\;k\neq i,$ (20)
where $\Gamma^{i}_{kj}$ are the spatial Christoffel symbols defined in Appendix A. We advance in time the contravariant positions and the orthonormal contravariant momenta, i.e. ${r}\equiv r^{i}\@vec{e}_{i}$ and ${u}\equiv u^{\hat{i}}\@vec{e}_{\hat{i}}$. To do so, we employ a generalization of the standard leapfrog method by Strang splitting the momentum equation (Strang 1968). Defining the positions centered in integer time steps and the ${u}$ velocities in half-time steps, the discretized equations of motion read:
${u}^{*}=\@vec{u}^{n-1/2}+\frac{\Delta t}{2}F_{L}^{n}\left(\@vec{r}^{n},\@vec{u}^{n-1/2}\right),$ (21)
${u}^{**}=\@vec{u}^{*}+\Delta tF_{G}^{n}\left(\@vec{r}^{n},\@vec{u}^{*}\right),$ (22)
${u}^{n+1/2}=\@vec{u}^{**}+\frac{\Delta t}{2}F_{L}^{n}\left(\@vec{r}^{n},\@vec{u}^{**}\right),$ (23)
${r}^{n+1}=\@vec{r}^{n}+\Delta tf^{n+1/2}\left(\@vec{r}^{n},\@vec{u}^{n+1/2}\right),$ (24)
@@
$F_{L}^{n}\left(\@vec{r}^{n},\@vec{u}^{n-1/2}\right)=\frac{\alpha\left(\@vec{r}^{n}\right)q}{m}\left(\@vec{E}^{n}\left(\@vec{r}^{n}\right)+\bar{\@vec{v}}^{n}\times\@vec{B}^{n}\left(\@vec{r}^{n}\right)\right),$ (25)
$F_{G}^{n}\left(\@vec{r}^{n},\@vec{u}^{*}\right)=\alpha\left(\@vec{r}^{n}\right)\Gamma\left(\@vec{u}^{*}\right)\@vec{g}\left(\@vec{r}^{n}\right)+\alpha\left(\@vec{r}^{n}\right)\@vec{H}\left(\@vec{r}^{n}\right)\cdot\@vec{u}^{*}+$
[…]
$f^{n+1/2}\left(\@vec{r}^{n},\@vec{u}^{n+1/2}\right)=\bar{\alpha}\@vec{u}^{n+1/2}-\bar{\@vec{\beta}},$ (27)
[…]
@@
In neutron stars, the general relativistic effects in the particle push may be seen as minor corrections to the particle motion very close to the star. In future works, we intend to detail the differences between simulations using a complete GR description and simulations that use a flat spacetime push with the frame-dragging effect in the field solver. This study may reveal if a numerically more expensive GR algorithm is necessary for neutron star magnetosphere simulations. Nevertheless, a complete implementation provides versatility to study plasma dynamics around more compact sources such as black holes.
2.5 Charge conservation
The only missing piece to complete the particle-in-cell algorithm is the evaluation of the charged current density in equation (16). This current term is the source term in the electric field evolution equation and captures the coupling between the electromagnetic fields and the charged particles. In the absence of sources, the Yee algorithm (Yee 1966) ensures that the time-independent Maxwell’s equations are always satisfied if initially satisfied. However, in the presence of sources, the only way to satisfy Gauss’s law (5) is through the charge conservation condition:
$\frac{\partial\rho}{\partial t}+\@vec{\nabla}\cdot\@vec{J}=0.$ (28)
An important challenge is then to design an algorithm capable of satisfying this condition to machine precision in an arbitrary curvilinear grid. Such algorithms are readily available for Cartesian grids for any interpolation order (Esirkepov 2001). The usual way to go around this challenge for curvilinear grids is to employ a divergence-cleaning algorithm, which is less efficient from a numerical point of view. Here, we present the general-relativistic generalization of the current deposition scheme first described in Cruz et al. (2023) that conserves charge to machine precision in the non-uniform grid in Fig. 1.

### v2

@@ 2.2 Grid and coordinates
Compact objects are naturally described in spherical coordinates. Using body-fitted coordinate grids simplifies the inclusion of system symmetries and boundary conditions. Another degree of freedom is the usage of linear or non-uniform grid spacings. A common feature in similar codes is to allow uniform grid spacings in $\log r$ (logarithmic coordinate) and $-\cos\theta$ (equal area coordinate). The first allows for higher resolution close to the central object, while the second enforces constant charge density of each macro-particle in the meridional direction (Belyaev, 2015a). These two properties have shown a reduction of numerical noise on the axis and increased code stability. Therefore, the OSIRIS-GR module has these two options available. Linear grid spacing options are also available, although these are not explored in the current paper. Another possibility available is to adopt logical coordinates, i.e. ($\tilde{r}$,$\tilde{\theta}$)=($\log r$,$-\cos\theta$), which possesses the two properties described above and retains uniform grid spacings.The background metric in the new coordinate system takes the form:
@@
$\tilde{A}^{\hat{r}}=\sqrt{{\gamma}_{\tilde{r}\tilde{r}}}\tilde{A}^{r}=\frac{\exp{\left(\tilde{r}\right)}}{\alpha}\exp{\left(-\tilde{r}\right)}A^{r}=\frac{1}{\alpha}A^{r}=\frac{1}{r}\frac{A^{\hat{r}}}{\alpha^{-1}}=A^{\hat{r}}.$ (4)
@@
Also, the OSIRIS-GR module introduces a new possibility towards modelling compact object magnetospheres via the half-domain setup. This simulation mode differs from the standard pole-to-pole meridional extension ($0\leq\theta\leq\pi$) by reducing the poloidal plane in half, thus possessing a pole-to-equator meridional extension ($20\leq\theta\leq\pi/$ 2). This novel numerical setup captures the global magnetospheric current closure while reducing the computational cost of emulating the complete poloidal plane. In this way, a higher number of macro-particles can be employed, increasing the statistical significance and reducing the associated numerical noise. This simulation mode is particularly interesting to study the pulsar polar cap dynamics, where the asymmetric mode development of the equatorial current sheet may be neglected to leading order.
@@ 2.3 Maxwell’s equations
In the 3+1 formalism, Maxwell’s equations take the form (Komissarov, 2011):
$\boldsymbol{\nabla}\cdot\boldsymbol{E}=4\pi\rho,$ (5)
$\boldsymbol{\nabla}\cdot\boldsymbol{B}=0,$ (6)
×$\boldsymbol{\nabla}\times\left(\alpha\boldsymbol{E}+\boldsymbol{\beta}\times\boldsymbol{B}\right)=-\frac{\partial\boldsymbol{B}}{\partial t},$ (7)
×$\boldsymbol{\nabla}\times\left(\alpha\boldsymbol{B}-\boldsymbol{\beta}\times\boldsymbol{E}\right)=\frac{\partial\boldsymbol{E}}{\partial t}+4\pi\left(\alpha\boldsymbol{j}-\rho\boldsymbol{\beta}\right),$ (8)
where $\boldsymbol{E}$, $\boldsymbol{B}$, $\boldsymbol{j}$ and $\rho$ are physical quantities measured by fiducial observers in co-rotation with the absolute space (ZAMOs). In spherical coordinates, Maxwell’s equations are better described in the integral form using the Kelvin–Stokes theorem on a Yee-lattice (Belyaev, 2015b; Cerutti et al., 2015). The OSIRIS-GR module adopts the spherical cell shown in Fig. 1.
@@
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{E}}\right)^{\hat{r}}_{i,j+1/2}=\frac{\alpha_{i}}{A^{\hat{r}}_{i,j+1/2}}\left(l^{\hat{\phi}}_{i,j+1}E^{\hat{\phi}}_{i,j+1}-l^{\hat{\phi}}_{i,j}E^{\hat{\phi}}_{i,j}\right),$ (9)
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{E}}\right)^{\hat{\theta}}_{i+1/2,j}=\frac{1}{A^{\hat{\theta}}_{i+1/2,j}}\left(l^{\hat{\phi}}_{i,j}\alpha_{i}E^{\hat{\phi}}_{i,j}-l^{\hat{\phi}}_{i+1,j}\alpha_{i+1}E^{\hat{\phi}}_{i+1,j}\right),$ (10)
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{E}}\right)^{\hat{\phi}}_{i+1/2,j+1/2}=\frac{l^{\hat{r}}_{i+1/2}}{A^{\hat{\phi}}_{i+1/2,j+1/2}}\left(\alpha_{i+1/2}\left(E^{\hat{r}}_{i+1/2,j}-E^{\hat{r}}_{i+1/2,j+1}\right)+\right.$
@@
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{B}}\right)^{\hat{r}}_{i+1/2,j}=\frac{\alpha_{i+1/2}}{A^{\hat{r}}_{i+1/2,j}}\left(l^{\hat{\phi}}_{i+1/2,j+1/2}B^{\hat{\phi}}_{i+1/2,j+1/2}-\right.$
@@
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{B}}\right)^{\hat{\theta}}_{i,j+1/2}=\frac{1}{A^{\hat{\theta}}_{i,j+1/2}}\left(l^{\hat{\phi}}_{i-1/2,j+1/2}\alpha_{i-1/2}B^{\hat{\phi}}_{i-1/2,j+1/2}-\right.$
@@
(×$\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{B}}\right)^{\hat{\phi}}_{i,j}=\frac{l^{\hat{r}}_{i}}{A^{\hat{\phi}}_{i,j}}\left(\alpha_{i}\left(B^{\hat{r}}_{i,j-1/2}-B^{\hat{r}}_{i,j+1/2}\right)+\right.$
@@
with $\boldsymbol{\tilde{E}}\equiv\alpha\boldsymbol{E}+\boldsymbol{\beta}\times\boldsymbol{B}$ and $\boldsymbol{\tilde{B}}\equiv\alpha\boldsymbol{B}-\boldsymbol{\beta}\times\boldsymbol{E}$, respectively. Expressions for the line and area elements are detailed in A. Up to this point, we have assumed $\beta^{\hat{\phi}}$ as the only non-zero component, with a purely radial dependence in $\alpha$, and neglected all azimuthal derivatives (axisymmetry). The evolution equations are then expressed as in the typical Yee (1966) scheme
$\boldsymbol{B}^{n+1/2}=\boldsymbol{B}^{n}-\frac{\Delta t}{2}\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{E}}\right)^{n},$ (15)
$\boldsymbol{E}^{n+1}=\boldsymbol{E}^{n}+\Delta t\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{B}}\right)^{n+1/2}-4\pi\Delta tJ^{n+1/2},$ (16)
$\boldsymbol{B}^{n+1}=\boldsymbol{B}^{n+1/2}-\frac{\Delta t}{2}\left(\boldsymbol{\nabla}\times\boldsymbol{\tilde{E}}\right)^{n+1},$ (17)
with $J\equiv\alpha\boldsymbol{j}-\rho\boldsymbol{\beta}$ and $n$ being the temporal index. Equation (16) also shows that to correctly push in time the electric field we need to gather the particle’s half-step charge density and current (see Sect. 2.5 and B). Due to the complex coupling introduced by equations (11) and (14), the poloidal components are pushed in time before the azimuthal ones in equations (16) and (17). By applying the Newtonian limit, we retrieve the Minkowski version of the evolution equations in Belyaev (2015b); Cerutti et al. (2015). Notice that (9)-(14) can be applied to any axisymmetric system given a diagonal spatial metric (e.g. spherical, cylindrical, toroidal).
@@ 2.4 Particle pusher
In the 3+1 formalism, the equations of motion for a charged particle (Thorne and MacDonald, 1982; Philippov et al., 2015a; Parfrey et al., 2019) are given by
$\frac{\mathrm{d}{r^{i}}}{\mathrm{d}t}\boldsymbol{e}_{i}=\frac{\alpha}{\Gamma}\boldsymbol{u}-\boldsymbol{\beta},$ (18)
$\frac{\mathrm{d}{u^{\hat{i}}}}{\mathrm{d}t}\boldsymbol{e}_{\hat{i}}=\frac{\alpha q}{m}\left(\boldsymbol{E}+\frac{\boldsymbol{u}}{\Gamma}\times\boldsymbol{B}\right)+\alpha\Gamma\boldsymbol{g}+\alpha\boldsymbol{H}\cdot\boldsymbol{u}+\boldsymbol{F}_{\mathrm{RR}}-\boldsymbol{F}_{\mathrm{Coord}},$ (19)
where $\boldsymbol{g}$ is the gravitational acceleration, $\boldsymbol{H}$ is the gravitomagnetic tensor, $\boldsymbol{p}=m\Gamma\boldsymbol{v}=m\boldsymbol{u}$ is the FIDO-measured momentum, $q$ and $m$ are the charge and mass of each macro-particle, $\Gamma=\sqrt{\varepsilon+\boldsymbol{u}\cdot\boldsymbol{u}}$ is the particle’s special relativistic Lorentz factor, and $\varepsilon=0\;\left(=1\right)$ for massless (massive) particles. Equation (18) shows that the particle velocity is composed of the FIDO-measured velocity, $\alpha\boldsymbol{u}/\Gamma$, and the FIDO’s velocity, $\boldsymbol{\beta}$. Equation (19) details which forces dictate the particle dynamics. From left to right, we identify the Lorentz, the gravitational, the gravitomagnetic, the classical radiation reaction (Vranic et al., 2016; Landau and Lifschitz, 1975), and the “coordinate” forces. The last term comprises the generalized fictitious forces associated with the use of the orthonormal vector basis:
$\boldsymbol{F}_{\mathrm{Coord}}=u^{\hat{i}}\frac{\mathrm{d}{\boldsymbol{e}_{\hat{i}}}}{\mathrm{d}t}=\frac{\sqrt{\gamma_{ii}}}{\sqrt{\gamma_{kk}}}\Gamma^{i}_{kj}u^{\hat{k}}\frac{\mathrm{d}x^{j}}{\mathrm{d}t}\boldsymbol{e}_{\hat{i}}=\Gamma^{\hat{i}}_{\hat{k}j}u^{\hat{k}}\frac{\mathrm{d}x^{j}}{\mathrm{d}t}\boldsymbol{e}_{\hat{i}},\;\;k\neq i,$ (20)
where $\Gamma^{i}_{kj}$ are the spatial Christoffel symbols defined in A. We advance in time the contravariant positions and the orthonormal contravariant momenta, i.e. $\boldsymbol{r}\equiv r^{i}\boldsymbol{e}_{i}$ and $\boldsymbol{u}\equiv u^{\hat{i}}\boldsymbol{e}_{\hat{i}}$. To do so, we employ a novel variant of the Strang splitting method to advance in time the equations of motion (Strang, 1968), which uses leapfrogged positions and velocities. Thus, defining the positions centered in integer time steps and the $\boldsymbol{u}$ velocities in half-time steps, the discretized equations of motion read:
$\boldsymbol{u}^{*}=\boldsymbol{u}^{n-1/2}+\frac{\Delta t}{2}F_{L}^{n}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{n-1/2}\right),$ (21)
$\boldsymbol{u}^{**}=\boldsymbol{u}^{*}+\Delta tF_{G}^{n}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{*}\right),$ (22)
$\boldsymbol{u}^{n+1/2}=\boldsymbol{u}^{**}+\frac{\Delta t}{2}F_{L}^{n}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{**}\right),$ (23)
$\boldsymbol{r}^{n+1}=\boldsymbol{r}^{n}+\Delta tf^{n+1/2}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{n+1/2}\right),$ (24)
@@
$F_{L}^{n}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{n-1/2}\right)=\frac{\alpha\left(\boldsymbol{r}^{n}\right)q}{m}\left(\boldsymbol{E}^{n}\left(\boldsymbol{r}^{n}\right)+\bar{\boldsymbol{v}}^{n}\times\boldsymbol{B}^{n}\left(\boldsymbol{r}^{n}\right)\right),$ (25)
$F_{G}^{n}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{*}\right)=\alpha\left(\boldsymbol{r}^{n}\right)\Gamma\left(\boldsymbol{u}^{*}\right)\boldsymbol{g}\left(\boldsymbol{r}^{n}\right)+\alpha\left(\boldsymbol{r}^{n}\right)\boldsymbol{H}\left(\boldsymbol{r}^{n}\right)\cdot\boldsymbol{u}^{*}+$
$\hskip 42.67912pt+\boldsymbol{F}_{RR}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{*}\right)-\boldsymbol{F}_{Coord}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{*}\right),$ (26)
$f^{n+1/2}\left(\boldsymbol{r}^{n},\boldsymbol{u}^{n+1/2}\right)=\bar{\alpha}\boldsymbol{u}^{n+1/2}-\bar{\boldsymbol{\beta}},$ (27)
where the terms with a bar highlight the approximated terms. The choice for $\bar{\boldsymbol{v}}$ defines if the simulation will be using the Boris (Boris and Shanny, 1972), the Vay (Vay, 2008), or the Higuera–Cary (HC) method (Higuera and Cary, 2017) for the electromagnetic force. For the systems we are interested in, the HC method is preferred as it is volume-preserving in the phase space and accurately captures the $E\times B$ drift even when under-resolving the cyclotron frequency (Higuera and Cary, 2017; Ripperda et al., 2018). The algorithm used to evaluate equations (22) and (24) is the second-order Heun’s (or Runge-Kutta) method.
@@
In neutron stars, the general relativistic effects in the particle push may be seen as minor corrections to the particle motion very close to the star. In future works, we intend to detail the differences between simulations using a complete GR description and simulations that use a flat spacetime push with the frame-dragging effect in the field solver (Philippov and Spitkovsky, 2018; Chen et al., 2020; Philippov et al., 2020; Bransgrove et al., 2022). This study may reveal if a numerically more expensive GR algorithm is necessary for neutron star magnetosphere simulations. Nevertheless, a complete implementation provides versatility to study plasma dynamics around more compact sources such as black holes.
2.5 Current deposition scheme
The only missing piece to complete the particle-in-cell algorithm is the evaluation of the charged current density in equation (16). This current term is the source term in the electric field evolution equation and captures the coupling between the electromagnetic fields and the charged particles. In the absence of sources, the Yee algorithm (Yee, 1966) ensures that the time-independent Maxwell’s equations are always satisfied if initially satisfied. However, in the presence of sources, the only way to satisfy Gauss’s law (5) is through the charge conservation condition:
$\frac{\partial\rho}{\partial t}+\boldsymbol{\nabla}\cdot\boldsymbol{J}=0.$ (28)
An important challenge is then to design an algorithm capable of satisfying this condition to machine precision in an arbitrary curvilinear grid. Such algorithms are readily available for Cartesian grids for any interpolation order (Esirkepov, 2001). The usual way to go around this challenge for curvilinear grids is to employ a divergence-cleaning algorithm; however, these algorithms are significantly less computationally efficient and, from a numerical point of view, their properties are less well known and explored (in a systematic way) . Here, we present the general-relativistic generalization of the current deposition scheme first described in Cruz et al. (2023) that conserves charge to machine precision in the non-uniform grid in Fig. 1, even for very high stellar compactness values. It is important to note that this general-relativistic charge conservation scheme is the first of its kind in the literature.